# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ Abstract
We study $t$-structures (on triangulated categories) that are closely related to weight structures. A $t$-structure couple $t=(\underline{C}_{t\leq 0},\underline{C}_{t\geq 0})$ is said to be adjacent to a weight structure $w=(\underline{C}_{w\leq 0},\underline{C}_{w\geq 0})$ if $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$. For a triangulated category $\underline{C}$ that satisfies the Brown representability property we prove that $t$ that is adjacent to $w$ exists if and only if $w$ is smashing (i.e., $\underline{C}_{w\geq 0}$ is $\coprod_{\underline{C}}$-closed). The heart ${\underline{Ht}}$ of this $t$ is the category of those functors ${\underline{Hw}}^{op}\to\operatorname{Ab}$ that respect products (here ${\underline{Hw}}$ is the heart of $w$); the result has important applications. We prove several more statements on constructing $t$-structures starting from weight structures; we look for a strictly orthogonal $t$-structure $t$ on some $\underline{C}^{\prime}$ (where $\underline{C},\underline{C}^{\prime}$ are triangulated subcategories of a common $\underline{D}$) such that $\underline{C}^{\prime}_{t\leq 0}$ (resp. $\underline{C}^{\prime}_{t\geq 0}$) is characterized by the vanishing of morphisms from $\underline{C}_{w\geq 1}$ (resp. $\underline{C}_{w\leq-1}$). Some of these results generalize properties of semi-orthogonal decompositions proved in the previous paper, and can be applied to various derived categories of (quasi)coherent sheaves on a scheme $X$ that is projective over an affine noetherian one. We also study hearts of orthogonal $t$-structures and their restrictions, and prove some statements on "reconstructing" weight structures from orthogonal $t$-structures.
The main tool of this paper is the notion of virtual $t$-truncations of (cohomological) functors; these are defined in terms of weight structures and "behave as if they come from $t$-truncations of representing objects" whether $t$ exists or not.
2020 Mathematics Subject Classification. 18G80, 14F08 (Primary) 18F20, 18G05, 18E10, 14A15, 14G40 (Secondary)
@@ Introduction
This paper is devoted to the study of those $t$-structures that are "closely related" to weight structures (on various triangulated categories).
Let us recall that $t$-structures on triangulated categories have become important tools in homological algebra since their introduction in [BBD82]. Respectively, their study and construction is an actual and non-trivial question. Next, in [Pau08] and [Bon10a] a rather similar notion of a weight structure $w$ on a triangulated category $\underline{C}$ was introduced. Moreover, in ibid. a $t$-structure $t=(\underline{C}_{t\leq 0},\underline{C}_{t\geq 0})$ was said to be adjacent to $w$ if $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$, and certain examples of adjacent structures were constructed. [^1] In contrast to ibid. and [Bon10b], in the current paper we use the so-called homological convention for $t$ and $w$, and say that $t$ is left adjacent to $w$ if $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$. Furthermore, in [Bon10b] for a $t$-structure $t$ on a triangulated category $\underline{C}^{\prime}$ that is related to $\underline{C}$ by means of a duality bi-functor a more general notion of orthogonality of a weight structure $w$ on $\underline{C}$ to $t$ was introduced.Also, the relationship between the hearts of adjacent and orthogonal structures was studied in detail.
Next, if $w$ is adjacent to $t$ then it determines $t$ uniquely and vice versa. Yet the only previously existing way of constructing $t$ if $w$ is given was to use certain "nice generators" of $w$ (see Definition 3.2.1(5), Theorem 3.2.2 of [Bon21], and of [Bon22a]). However, already in [Bon10a] the notion of virtual $t$-truncations for (co)homological functors was introduced, and it was proved that virtual $t$-truncations possess several nice properties. In particular, it was demonstrated that these are closely related to $t$-structures (whence the name) even though they are defined in terms of weight structures only.
In the current paper we propose a new construction method. We prove that adjacent and orthogonal $t$-structures can be constructed using virtual $t$-truncations whenever certain "Brown representability-type" assumptions on $\underline{C}$ (and $\underline{C}^{\prime}$) are known. Respectively, our results yield the existence of some new families of $t$-structures.
@@
Let $\underline{C}$ be a smashing triangulated category that satisfies the following Brown representability property: any functor $\underline{C}^{op}\to\operatorname{Ab}$ that respects ($\underline{C}^{op}$)-products is representable.
@@
Note here that (smashing) triangulated categories satisfying the Brown representability property are currently really popular in several areas of mathematics (thanks to the foundational results of A. Neeman and others); in particular, this property holds if either $\underline{C}$ or $\underline{C}^{op}$ is compactly generated. Moreover, it is easy to construct vast families of smashing weight structures on $\underline{C}$ (at least) if $\underline{C}$ is compactly generated; see Remark 3.2.4(1) below.
@@
Certainly the dual to Theorem 0.1 is valid as well. Moreover, if $\underline{C}^{op}$ satisfies the dual Brown representability property and $w$ is both cosmashing and smashing then the right adjacent $t$-structure $t$ (i.e., $\underline{C}_{t\leq 0}=\underline{C}_{w\leq 0}$) restricts to the subcategory of compact objects of $\underline{C}$ as well as to all other "levels of smallness" for objects (see Theorem 3.4.1(3)). Combining this statement with an existence of weight structures Theorem 3.1 of [KeN13] we obtain a statement on $t$-structures extending Theorem 7.1 of ibid.
@@
We also prove some alternative versions of Theorem 0.1 that can be applied to "quite small" triangulated categories. Instead of the Brown representability condition for $\underline{C}$ one can demand it to satisfy the $R$-saturatedness one instead (see Definition 4.2.1(2) below; this is an "$R$-linear finite" version of the Brown representability). Then for any bounded $w$ on $\underline{C}$ there will exist a $t$-structure adjacent to it. Moreover, this statement also holds for essentially bounded weight structures; those generalize both bounded weight structures and semi-orthogonal decompositions. According to a saturatedness statement from [Nee18] this result can be applied to the derived category $D^{perf}(X)$ of perfect complexes on a regular scheme that is proper over the spectrum of a Noetherian ring $R$ (see Corollary 4.2.2(3)). We also study restrictions of orthogonal $t$-structures to subcategories corresponding to certain bounds and support conditions; see Propositions 2.3.5 and 3.3.4, Corollaries 4.2.2(2) and 4.2.5, and Theorem 4.2.4. We demonstrate that this result can be applied to certain derived categories of (quasi)coherent sheaves; this includes the case of a singular $X$ (see Propositions 3.3.5 and 4.2.6).
@@
Lastly, we try to prove some converse of these results. So we study the question when a (fixed) $t$-structure $t$ on a triangulated category $\underline{C}$ is orthogonal to a weight structure $w$ on a "large enough" subcategory $\underline{C}^{\prime}\subset\underline{C}$. Roughly, this is the case if and only if the heart ${\underline{Ht}}$ has enough projectives and those "lift" to ${\underline{Hw}}$ (and the latter condition follows from duals to certain Brown representability-type conditions). However, it appears to be difficult to find the conditions that ensure the existence of an adjacent weight structure (so, it is on $\underline{C}$ itself). We prove some statements of this sort; yet these results are more difficult to apply than the aforementioned "converse" ones, and the methods of their proofs are less original.
@@
The author certainly does not claim that the methods of the current paper are the most general among the existing methods for constructing $t$-structures. In particular, if $\underline{C}$ is a well generated triangulated category (categories of this sort are "very common") then the recent Theorem 2.3 of [Nee21] gives all those $t$-structures that are generated by sets of objects of $\underline{C}$ in the sense of Definition 3.2.1(5) below.
Now, if $\underline{C}$ is well generated then is generated by a set of its objects (as its own localizing subcategory; see Definition 3.2.1(3) below). Thus for any smashing weight structure $w$ on it Proposition 2.3.2(10) of [Bon21] essentially says the following: there exists a set $\mathcal{P}\subset\underline{C}_{w\geq 0}$ such that the class $\underline{C}_{w\geq 0}$ is the smallest cocomplete pre-aisle that is "generated" by $\mathcal{P}$ in the sense of [Nee21, §0] (cf. Discussion 1.16 of ibid.). Hence Theorem 2.3 of [Nee21] says that there exists a $t$-structure on $\underline{C}$ such that $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$ (cf. Remark 1.1.6(2) below). Thus loc. cit. generalizes the existence of $t$ part of our Theorem 0.1.
@@
In §1 we give some definitions and conventions, and recall some basics on $t$-structures, weight structures, and semi-orthogonal decompositions.
@@
In §4 we discuss certain bounded versions of the results of the previous section. Our main statement concern the existence of adjacent and orthogonal $t$-structures in $R$-saturated triangulated categories and the restriction of orthogonal $t$-structures to subcategories corresponding to certain bounds and support conditions. We apply our general statements to various derived categories of (quasi)coherent sheaves. To formulate our statements in a more general case we recall the notions of essentially bounded (above, below, or both) objects and weight structures. In particular, we are able to generalize the central results of [Bon24].
In §5 we try to answer the question which $t$-structures are adjacent to weight structures. In certain situations we prove that a weight structure adjacent to a given $t$ exists whenever ${\underline{Ht}}$ has enough projectives and certain additional assumptions are fulfilled.
@@
Remark 0.3.
1. The current text is a certain modification of the preprint [Bon19]. The main idea is to avoid the somewhat ad hoc (and restrictive) notion of reflection of categories (see Definition 2.2.1 of ibid.); cf. Remark 2.4.5 below. Moreover, several new general results were added; this includes the theory of coproductive extensions, (most of) §4, §5.3, and subcategories corresponding to certain support conditions (see Propositions 2.3.5, 3.3.4, 3.3.5, and 4.2.6(2), and Corollary 4.2.2(2)).
2. Since [Bon19] was already cited in some papers, it makes sense to note the following: Theorem 3.2.3(I) of ibid. coincides with Theorem 0.1 of the current paper, Theorem 3.3.1 of ibid. almost coincides with Theorem 3.3.2 below, Theorem 5.3.1(I.1) of ibid. is contained in Proposition 5.2.1, Theorem 5.3.1(IV) of ibid. was generalized to Theorem 5.3.3, and strongly ${\aleph_{0}}$–well generated weight structures (that were treated in §3.3 of [Bon19]) are discussed in §3.3 (yet see Remark 3.3.3(1) below).
3. The work on [Bon19] inspired the author to write [Bon24]. We note that semi-orthogonal decompositions (that are the main subject of ibid.) are particular case of weight structures; see Proposition 1.2.6 below for more detail. Next, both in [Bon24] and in (the justificatio of) Remark 4.2.7(1) below the main idea is to extend a weight structure to a bigger (smashing) category and to pass to an orthogonal $t$-structure (see Definitions 1.2.2(4) and 1.3.1 below). It is no wonder that applying these methods in the setting of (general) weight structures is somewhat more difficult than doing this for semi-orthogonal decompositions. On the other hand, these difficulties make the current paper somewhat more interesting and also "more original".
@@ 1 A reminder on $t$-structures and weight structures
In §1.2 we recall some of the theory of weight structures and semi-orthogonal decompositions.
@@ 1.1 Some categorical and $t$-structure notation
  • A class $\mathcal{P}\subset\operatorname{Obj}\underline{C}$ is said to be extension-closed if it is closed with respect to extensions and contains 00.
@@
  • We will say that $\underline{A}^{0}\subset\underline{A}$ is an abelian subcategory of $\underline{A}$ if $\underline{A}$ is its (strictly full) subcategory that contains the $\underline{A}$-kernel and the $\underline{A}$-cokernel of any $\underline{A}^{0}$-morphism.
@@
  • We will say that an additive covariant (resp. contravariant) functor from $\underline{C}$ into $\underline{A}$ is homological (resp. cohomological) if it converts distinguished triangles into long exact sequences. For a (co)homological functor $H$ and $i\in{\mathbb{Z}}$ we will write $H_{i}$ (resp. $H^{i}$) for the composition $H\circ[-i]$.
Remark 1.1.
Obviously, if $\underline{C}_{1},\underline{C}_{2}$ are triangulated subcategories of $\underline{C}$ then $\underline{C}_{1}\cap\underline{C}_{2}$ is its triangulated subcategory as well; recall the strictness assumption.
Similarly, if $\underline{A}_{1},\underline{A}_{2}$ are weak Serre subcategory of $\underline{A}$ then $\underline{A}_{1}\cap\underline{A}_{2}$ is its weak Serre subcategory as well.
To generalize our formulations and relate them to the results of [Bon10b] we recall Definition 2.5.1(1) of ibid.
Definition 1.1.1.
The following example of a duality will be the main one throughout this paper.
Lemma 1.1.2.
Assume $\underline{C},\underline{C}^{\prime}\subset\underline{D}$. Then the restriction of the bi-functor $\underline{D}(-,-)$ to $\underline{C}^{op}\times\underline{C}^{\prime}$ gives a duality $\underline{C}^{op}\times\underline{C}^{\prime}\to\operatorname{Ab}$.
Proof.
This statement is obvious; it also easily follows from Proposition 2.5.6(1) of [Bon10b]. ∎
@@
    $\underline{C}_{t\ ge0}$ is closed with respect to all (small) coproducts that exist in $\underline{C}$.
@@
2. Even though in [BBD82] where $t$-structures were introduced and in several preceding papers of the author the "cohomological convention" for $t$-structures was used, in the current text we use the homological convention; the reason for this is that it is coherent with the homological convention for weight structures (see Remark 1.2.3(3) below). Respectively, our notation $\underline{C}_{t\geq 0}$ corresponds to the class $\underline{C}^{t\leq 0}$ in the cohomological convention.
@@
Conversely, to check whether $t$ restricts to ${\underline{C}^{0}}$ it clearly suffices to verify that any object $M$ of ${\underline{C}^{0}}$ possesses a decomposition of the sort (1.1.1) inside ${\underline{C}^{0}}$. Now, if $L_{t}(M)\in\operatorname{Obj}{\underline{C}^{0}}$ then $R_{t}(M)$ is an object of ${\underline{C}^{0}}$ as well since ${\underline{C}^{0}}$ is a triangulated subcategory of $\underline{C}$.
2. $\underline{C}^{2}$ is triangulated since $\underline{C}^{0}$ and $\underline{C}^{1}$ are (fully) strict subcategories of $\underline{C}$, whereas different choices of cones of a $\underline{C}$-morphisms are connected by (non-unique) isomorphisms.
The existence of the restricted weight structure statement is easy as well; one can immediately deduce this statement from assertion 1. Lastly, the heart assertion is an immediate consequence of our definitions. ∎

### v2

@@ Abstract
We study $t$-structures (on triangulated categories) that are closely related to weight structures. A $t$-structure couple $t=(\underline{C}_{t\leq 0},\underline{C}_{t\geq 0})$ is said to be adjacent to a weight structure $w=(\underline{C}_{w\leq 0},\underline{C}_{w\geq 0})$ if $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$. For a triangulated category $\underline{C}$ that satisfies the Brown representability property we prove that $t$ that is adjacent to $w$ exists if and only if $w$ is smashing (i.e., $\underline{C}_{w\geq 0}$ is $\coprod_{\underline{C}}$-closed); the heart of this $t$ is the category of those functors ${\underline{Hw}}^{op}\to\operatorname{Ab}$ that respect products (here ${\underline{Hw}}$ is the heart of $w$). This result has important applications.
We prove several more statements on constructing (orthogonal) $t$-structures starting from weight structures. Some of these results generalize properties of semi-orthogonal decompositions proved in an earlier paper, and can be applied to various derived categories of (quasi)coherent sheaves on a scheme $X$ that is projective over an affine noetherian one. We also study hearts of orthogonal $t$-structures and their restrictions, and prove some statements on ’reconstructing’ weight structures from orthogonal $t$-structures.
The main tool of this paper is the notion of virtual $t$-truncations of (cohomological) functors; these are defined in terms of weight structures and behave as if they come from $t$-truncations of representing objects’ whether $t$ exists or not.
2020 Mathematics Subject Classification. 18G80, 14F08 (Primary) 18F20, 18G05, 18E10, 14A15 (Secondary)
@@ Introduction
This paper is devoted to the study of those $t$-structures that are closely related to weight structures (on various triangulated categories).
Let us recall that $t$-structures on triangulated categories have become important tools in homological algebra since their introduction in [BBD82]. Respectively, their study and construction is an actual and non-trivial question. Next, in [Pau08] and [Bon10a] a rather similar notion of a weight structure $w$ on a triangulated category $\underline{C}$ was introduced. Moreover, in ibid. a $t$-structure $t=(\underline{C}_{t\leq 0},\underline{C}_{t\geq 0})$ was said to be (left) adjacent to $w$ if $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$, and certain examples of adjacent structures were constructed. Furthermore, in [Bon10b] for a $t$-structure $t$ on a triangulated category $\underline{C}^{\prime}$ that is related to $\underline{C}$ by means of a duality bi-functor a more general notion of orthogonality of a weight structure $w$ on $\underline{C}$ to $t$ was introduced. Also, the relationship between the hearts of adjacent and orthogonal structures was studied in detail.
Next, if $w$ is adjacent to $t$ then it determines $t$ uniquely and vice versa. Yet the only previously existing way of constructing $t$ if $w$ is given was to use certain ’nice generators’ of $w$ (see Definition 3.2.1(5), Theorem 3.2.2 of [Bon21], and Theorem 2.4.2(II.2) of [Bon22]). However, already in [Bon10a] the notion of virtual $t$-truncations for (co)homological functors was introduced, and it was proved that virtual $t$-truncations possess several nice properties. In particular, it was demonstrated that these are closely related to $t$-structures (whence the name) even though they are defined in terms of weight structures only.
In the current paper we propose a new construction method. We prove that adjacent and orthogonal $t$-structures can be constructed using virtual $t$-truncations whenever certain Brown representability-type assumptions on $\underline{C}$ (and $\underline{C}^{\prime}$) are known. Respectively, our results yield the existence of some new families of $t$-structures.
@@
Let $\underline{C}$ be a smashing triangulated category that satisfies the following Brown representability property: any functor $\underline{C}^{op}\to\operatorname{Ab}$ that respects ($\underline{C}^{op}$)-products is representable. [^1] Equivalently, a functor is representable if and only if it converts $\underline{C}$-coproducts into products of abelian groups.
@@
Note here that (smashing) triangulated categories satisfying the Brown representability property are currently really popular in several areas of mathematics (thanks to the foundational results of A. Neeman, H. Krause and others); in particular, this property holds if either $\underline{C}$ or $\underline{C}^{op}$ is compactly generated. Moreover, it is easy to construct vast families of smashing weight structures on $\underline{C}$ (at least) if $\underline{C}$ is compactly generated; see Remark 3.2.4(1) below.
@@
Certainly the dual to Theorem 0.1 is valid as well. Moreover, if $\underline{C}^{op}$ satisfies the dual Brown representability property and $w$ is both cosmashing and smashing then its right adjacent $t$-structure $t$ (characterized by the assumption $\underline{C}_{t\leq 0}=\underline{C}_{w\leq 0}$) restricts to the subcategory of compact objects of $\underline{C}$ as well as to all other ’levels of smallness’ for objects (see Theorem 3.4.1(3)). Combining this statement with an existence of weight structures Theorem 3.1 of [KeN13] we obtain a statement on $t$-structures extending Theorem 7.1 of ibid.
@@
We also prove some alternative versions of Theorem 0.1 that can be applied to ’quite small’ triangulated categories. Instead of the Brown representability condition for $\underline{C}$ one can demand it to satisfy the $R$-saturatedness one (see Definition 4.2.1(2) below; this is an $R$-linear finite analogue of the Brown representability). Then for any bounded $w$ on $\underline{C}$ (that is, $\cup_{i\in{\mathbb{Z}}}\underline{C}_{w\leq 0}[i]=\cup_{i\in{\mathbb{Z}}}\underline{C}_{w\geq 0}[i]$) there will exist a $t$-structure adjacent to it. Moreover, this statement also holds for essentially bounded weight structures (see Corollary 4.2.2(1)); those generalize both bounded weight structures and semi-orthogonal decompositions. According to a saturatedness statement from [Nee18] this result can be applied to the derived category $D^{perf}(X)$ of perfect complexes on a regular scheme that is proper over the spectrum of a Noetherian ring $R$ (see Corollary 4.2.2(3)). We also study restrictions of orthogonal $t$-structures to subcategories corresponding to certain bounds and support conditions; see Propositions 2.3.5 and 3.3.4, Corollaries 4.2.2(2) and 4.2.5, and Theorem 4.2.4. We demonstrate that these results can be applied to certain derived categories of (quasi)coherent sheaves; this includes the case of a singular $X$ (that is proper over $\operatorname{Spec}R$; see Propositions 3.3.5 and 4.2.6).
@@
Lastly, we try to prove some converse of these results. So we study the question when a (fixed) $t$-structure $t$ on a triangulated category $\underline{C}$ is orthogonal to a weight structure $w$ on a ’large enough’ subcategory $\underline{C}^{\prime}\subset\underline{C}$. Roughly, this is the case if and only if the heart ${\underline{Ht}}$ has enough projectives and those lift to $\underline{C}$ in a certain way (and the latter condition follows from duals to certain Brown representability-type conditions). However, it appears to be difficult to find the conditions that ensure the existence of an adjacent weight structure (so, $\underline{C}^{\prime}=\underline{C}$). We prove some statements of this sort; yet these results are more difficult to apply than the aforementioned converse ones, and the methods of their proofs are less original.
@@
1. The author certainly does not claim that the methods of the current paper are the most general among the existing methods for constructing $t$-structures. Yet they give more information on the hearts of the corresponding $t$-structures than some other results in this direction; cf. Remark 6.4 below for more detail.
2. The current text is a certain modification of the preprint [Bon19]. We will say more on the comparison between these two texts in Remark 6.3 below.
3. The work on [Bon19] inspired the author to write [Bon24b]. We note that semi-orthogonal decompositions (that are the main subject of ibid.) are particular (and very important!) case of weight structures; see Proposition 1.2.6 below for more detail. Next, both in [Bon24b] and in §4.2 below the main idea is to extend a weight structure to a bigger (smashing) category and to pass to an orthogonal $t$-structure (see Definitions 1.2.2(4) and 1.3.1 below). It is no wonder that applying these methods in the setting of (general) weight structures is somewhat more difficult than doing this for semi-orthogonal decompositions. On the other hand, these difficulties make the current paper somewhat more interesting and also ’more original’.
@@
In §1 we give some definitions and conventions, and recall some basics on $t$-structures, weight structures, and semi-orthogonal decompositions. The most original statement of this section is Proposition 1.2.6 that states that semi-orthogonal decompositions are precisely those weight structures whose heart is zero; these weight structures are also characterized by the uniqueness of the corresponding decompositions.
@@
In §4 we discuss certain bounded versions of the results of the previous section. Our main statement concern the existence of adjacent and orthogonal $t$-structures in $R$-saturated triangulated categories and the restriction of orthogonal $t$-structures to subcategories corresponding to certain bounds and support conditions. We apply our general statements to various derived categories of (quasi)coherent sheaves. To formulate our statements in a more general setting we recall the notions of essentially bounded (above, below, or both) objects and weight structures. In particular, we are able to generalize the central results of [Bon24b].
In §5 we address the question which $t$-structures are adjacent to weight structures. In certain situations we prove that a weight structure adjacent to a given $t$ exists whenever ${\underline{Ht}}$ has enough projectives and certain additional assumptions are fulfilled.
§6 contains some additional remarks that relate this paper with some earlier ones. None of them appear to be really important.
@@ 1 A reminder on $t$-structures and weight structures
In §1.2 we recall some of the theory of weight structures and semi-orthogonal decompositions. In particular, Proposition 1.2.6 states that semi-orthogonal decompositions are precisely those weight structures whose heart is zero.
@@ 1.1 Some categorical and $t$-structure notation
  • A class $\mathcal{P}\subset\operatorname{Obj}\underline{C}$ is said to be extension-closed if $0\in\mathcal{P}$ and $\mathcal{P}\star\mathcal{P}\subset\mathcal{P}$.
@@
    We will say that $\mathcal{P}$ densely generates $\underline{C}$ if the envelope of $\cup_{i\in{\mathbb{Z}}}\mathcal{P}[i]$ equals $\underline{C}$.
@@
  • We will say that an additive covariant (resp. contravariant) functor from $\underline{C}$ into $\underline{A}$ is homological (resp. cohomological) if it converts distinguished triangles into long exact sequences.
    For a (co)homological functor $H$ and $i\in{\mathbb{Z}}$ we will write $H_{i}$ (resp. $H^{i}$) for the composition $H\circ[-i]$.
@@
    $\underline{C}_{t\geq 0}$ is closed with respect to all (small) coproducts that exist in $\underline{C}$.
@@
2. Even though in [BBD82] where $t$-structures were introduced and in several preceding papers of the author the so-called cohomological convention for $t$-structures was used, in the current text we use the homological convention; the reason for this is that it is coherent with the homological convention for weight structures (see Remark 1.2.3(3) below). Respectively, our notation $\underline{C}_{t\geq 0}$ corresponds to the class $\underline{C}^{t\leq 0}$ in the cohomological convention.
@@
Conversely, to check whether $t$ restricts to ${\underline{C}^{0}}$ it clearly suffices to verify that any object $M$ of ${\underline{C}^{0}}$ possesses a decomposition of the sort (1.1.1) inside ${\underline{C}^{0}}$. Now, if $L_{t}(M)\in\operatorname{Obj}{\underline{C}^{0}}$ then $R_{t}(M)$ is an object of ${\underline{C}^{0}}$ as well, since ${\underline{C}^{0}}$ is a triangulated subcategory of $\underline{C}$.
2. $\underline{C}^{2}$ is triangulated since $\underline{C}^{0}$ and $\underline{C}^{1}$ are (fully) strict triangulated subcategories of $\underline{C}$, whereas different choices of cones of a $\underline{C}$-morphisms are connected by (non-unique) isomorphisms.
The existence of the restricted $t$-structure statement is easy as well; one can immediately deduce this statement from assertion 1. Lastly, the heart assertion is an immediate consequence of our definitions. ∎